OpenStudy (anonymous):

Samsonite16 PLEASE HELP AND WILL GIVE MEDAL! Simplify and select the answer with the appropriate restrictions for the variables. 6x^2-3x/3x^2

OpenStudy (anonymous):

factor out 3 from 6x^2 -3x . What do you get?

OpenStudy (anonymous):

? lol

OpenStudy (anonymous):

factored out you'd get 3x(2x-1)/3x^2 Again anything divided by 0 is undefined x cannot be 0

OpenStudy (anonymous):

thx

OpenStudy (aum):

(6x^2 - 3x) / 3x^2 = 3x(2x - 1) / 3x^2 = (2x - 1) / x = 2 - 1/x x cannot be zero.

OpenStudy (anonymous):

thank you @aum and @study100 :)

OpenStudy (aum):

You are welcome.